# PRIMA.CPP: Speeding Up 70B-Scale LLM Inference on Low-Resource Everyday Home Clusters

### v1

@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
For the device loading time, processing cores must load data from RAM/VRAM into registers before executing instructions, which incurs latency. However, the theoretical bandwidth of memory cards cannot estimate this latency because apps often fail to fully utilize the bandwidth, and multi-level caching also has a significant influence. To capture these effects, our device profiler implements an operator to read data from RAM/VRAM into registers. By measuring its latency with data volumes similar to the tensor sizes, we have the practical throughput, denoted as $\{\mathcal{T}_{m}^{\text{cpu}},\mathcal{T}_{m}^{\text{cuda}},\mathcal{T}_{m}^{\text{metal}}\}$. Next, we count the data volume that needs to be loaded into registers during each token step, which typically consists of the weight data and the key-value cache. In the model profiler, we record the total bytes of weight data for the input and output layer as $b_{i},b_{o}$, and for each layer as $b$. Additionally, the key-value cache size for each layer is $2(h_{k}e_{k}+h_{v}e_{v})n_{kv}$, where $n_{kv}$ is the number of tokens for which the cache is stored. Then, the device loading time $T_{m}^{\text{dev\_load}}$ can be expressed as $T_{m}^{\text{dev\_load}}=(\frac{l_{m}-l_{m}^{\text{gpu}}}{\mathcal{T}_{m}^{\text{cpu}}}+\frac{l_{m}^{\text{gpu}}}{\mathcal{T}_{m}^{\text{gpu}}})(b+2(h_{k}e_{k}+h_{v}e_{v})n_{kv})+\frac{b_{i}/V+b_{o}}{\mathcal{T}_{m}^{\text{cpu}}}\cdot\mathbb{I}_{m=1}$, where $\mathcal{T}_{m}^{\text{gpu}}$ depends on the hardware: it equals to $\mathcal{T}_{m}^{\text{metal}}$ if the GPU uses Metal, or $\mathcal{T}_{m}^{\text{cuda}}$ for CUDA, and $V$ is the vocabulary size.
@@
Estimation of disk loading latency $T_{m}^{\text{disk}}$. Prima.cpp is designed to run on memory-constrained home devices, so it cannot load the entire model weights into physical RAM. To address this, prima.cpp uses mmap to manage model weights. By using mmap, model weights are loaded into memory from disk only when needed for computation, and the OS will release inactive mmap-ed pages when memory pressure is high. This prevents OOM issues but incurs significant disk I/O latency because mmap must reload the model weights when they are accessed again after being released. To estimate this disk loading latency, it is necessary to determine the data volume that mmap needs to load in each token step. This is a challenging task because different OSs have very different memory management behaviors and great dynamics.
On macOS (without Metal) and Linux, the OS gradually reclaims memory. When memory pressure is moderate, i.e., when $b_{lio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}>d_{m}^{\text{avail}}$, mmap-ed pages are released incrementally until the pressure is alleviated. As a result, some weight data remain in the page cache, and the amount of data that mmap needs to reload is $\max(b_{lio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}},0)$, where $b_{lio}=(l_{m}-l_{m}^{\text{gpu}})b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}$, $2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})$ is the key-value cache size, and $c^{\text{cpu}}$ is the total computing buffer size. If CUDA is enabled on Linux, the model weights in private VRAM are locked by the CUDA driver, keeping them resident so no disk I/O occurs. Therefore, the disk loading latency for macOS (without Metal) and Linux can be estimated as
@@
For $T_{m,\text{macOS(no Metal)}}^{\text{disk}}$, $l_{m}^{\text{gpu}}=0$ and $s_{m}^{\text{disk}}$ is the random read throughput. On Linux, mmap is configured for sequential access, so $s_{m}^{\text{disk}}$ is now the sequential read throughput.
However, when Metal is enabled on macOS, the behavior changes. Metal loads mmap pages into the shared memory, and the OS prioritizes retaining these pages. That is, the OS is more inclined to swap out or compress active pages while keeping mmap-ed model weight pages in shared memory intact. However, when memory is exhausted (with free and inactive pages exhausted, the compression pool nearing saturation, and heavy swap usage), macOS will release these mmap-ed pages in a more aggressive manner. This may cause the entire model weights to be repeatedly loaded and released. As a result, when the required memory exceeds the total available memory, i.e., when $l_{m}b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}l_{m}+c^{\text{cpu}}+c^{\text{gpu}}>d_{m,\text{metal}}^{\text{avail}}$, device $d_{m}$ needs to reload $l_{m}b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}$ bytes in each token step. Here, $d_{m,\text{metal}}^{\text{avail}}$ denotes the maximum working set size recommended by Metal. By measuring the random read throughput $s_{m}^{\text{disk}}$ of disk, we can calculate the disk loading latency for macOS (with Metal) as:
@@
$+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}l_{m}+c^{\text{cpu}}+c^{\text{gpu}}-d_{m,\text{metal}}^{\text{avail}}\Big{)},b_{i}/V\Big{)}.$ (16)
When running on Android devices, the OS prioritizes swapping out inactive pages to disk, such as memory used by background apps, to ensure that the active app runs smoothly. As a result, the available RAM for prima.cpp can be higher than initially expected, as the OS will swap out cold pages to disk, freeing up additional space in memory. Thus, on Android, the number of bytes that mmap needs to reload is $\max(b_{bio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}}-d_{m}^{\text{swapout}},0)$, where $d_{m}^{\text{swapout}}=\min(\max(0,b_{bio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}}),\min(d_{m}^{\text{bytes\_can\_swap}},d_{m}^{\text{swap\_avail}}))$ represents the data bytes that are swapped out to disk, $d_{m}^{\text{bytes\_can\_swap}}$ is the data bytes of currently used memory that can be swapped out, and $d_{m}^{\text{swap\_avail}}$ is the total available swap space on the device. Then we have:
@@
By aggregating them, we obtain a unified expression compatible for cross-platform devices:
$T_{m}^{\text{disk}}=$ $T_{m,\text{macOS (no Metal)}}^{\text{disk}}\cdot\mathbb{I}_{\text{macOS (no Metal)}}+T_{m,\text{macOS (with Metal)}}^{\text{disk}}\cdot\mathbb{I}_{\text{macOS (with Metal)}}$ (18)
@@
where $\mathbb{I}_{\text{macOS (no Metal)}},\mathbb{I}_{\text{macOS (with Metal)}},\mathbb{I}_{\text{Linux}},\mathbb{I}_{\text{Android}}$ are indicator functions. This expression can be easily extended to include new OSs, e.g., Windows will be added in future updates.
Estimation of network communication latency $T_{m}^{\text{comm}}$. In prima.cpp, devices are logically interconnected in a ring structure, where each device receives input from its predecessor, processes a layer window, and sends the output to its successor. After a device completes the computation for one layer window, it transmits the result ($e$ values in F32 format, totaling $4e$ bytes) to the next device for further computation on the next layer window. Therefore, during a token step, the number of network communications on device $d_{m}$ equals the number of layer windows, which is $\mathcal{W}_{m}=\left\lfloor\frac{L}{W}\right\rfloor+\min(1,\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$. By measuring the latency $t_{m}^{\text{comm}}$ for transmitting $4e$ bytes between adjacent devices, we can estimate the network communication latency on device $d_{m}$ as:
@@
By aggregating all these latencies, we have the objective as:
@@
$+\frac{\max(l_{m}b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}l_{m}+c^{\text{cpu}}-d^{\text{avail}}_{m},0)}{s_{m}^{\text{disk}}}\cdot\mathbb{I}_{\text{macOS (no Metal)}}$
@@
$\qquad\qquad b_{i}/V\Big{)}\cdot\mathbb{I}_{\text{macOS (with Metal)}}$
@@
To remove the max operator, we decompose the disk loading latency into multiple terms, separately accounting for cases where memory is sufficient or insufficient. Let $\mathcal{M}$ be the set of all devices, and $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$ be the subsets of devices that satisfy the respective conditions in Cases 1-4, where $\mathcal{M}_{1}\cap\mathcal{M}_{2}\cap\mathcal{M}_{3}\cap\mathcal{M}_{4}=\emptyset$ and $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}\cup\mathcal{M}_{4}=\mathcal{M}$.
@@
Case 4 (OS with sufficient RAM or low disk speed): In these cases, the physical RAM is large enough to hold the model weights or the disk speed is too slow (i.e., $s_{m}^{\text{disk}}<s_{\text{threshold}}^{\text{disk}}$). As a result, no disk loading is expected, except for the latency incurred during lookup table access, thus $T_{m}^{\text{disk}}=\frac{b_{i}}{s_{m}^{\text{disk}}V},m\in\mathcal{M}_{4}$.
@@
Let $\frac{L}{W}$ be an integer, i.e., $R=0$, where $W=\sum_{m\in\mathcal{M}}w_{m}$, all devices are assigned an equal number of windows and all the windows are filled.
Now, we have $l_{m}=\frac{w_{m}L}{W}$, $l_{m}^{\text{gpu}}=\frac{n_{m}L}{W}$, $\mathcal{W}_{m}=\frac{L}{W}$. Let $b^{\prime}=b+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}$, $\alpha_{m}=\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+t_{m}^{\text{kv\_cpy,cpu}}+\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{cpu}}}$, $\beta_{m}=\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{gpu},q}}-\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+t_{m}^{\text{kv\_cpy,gpu}}-t_{m}^{\text{kv\_cpy,cpu}}+\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{gpu}}}-\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{cpu}}}$, $\xi_{m}=(t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}})(1-\mathbb{I}_{m}^{\text{UMA}})+t_{m}^{\text{comm}}$, $\kappa=\sum_{q\in\mathcal{Q}}\frac{f_{1,\text{out}}^{q}}{s_{1}^{\text{cpu},q}}+\frac{b_{i}/V+b_{o}}{\mathcal{T}_{1}^{\text{cpu}}}+\frac{b_{i}/V}{s_{1}^{\text{disk}}}+\frac{b_{o}}{s_{1}^{\text{disk}}}\cdot\mathbb{I}_{1\notin\mathcal{M}_{4}}+\sum_{m\in\mathcal{M}_{1}\cup\mathcal{M}_{3}}\frac{c^{\text{cpu}}-d_{m}^{\text{avail}}-d_{m}^{\text{swapout}}\cdot\mathbb{I}_{\text{Android}}}{s_{m}^{\text{disk}}}$, where $\alpha_{m},\beta_{m},\xi_{m}$ are platform-specific constants and $\kappa$ is a global constant. Then, we add the first general term to the three platform-specific terms and obtain:
@@
This objective is a sum over three sets, $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$. Each summand involves expressions linear in $w_{m}$ and $n_{m}$, plus constant terms specific to the platform. To clarify the form, we define a linear function $f(a,b,c)=aw_{m}+bn_{m}+c$, where the platform-specific constants $a,b,c$ are independent of the decision variables $w_{m}$ and $n_{m}$. Consequently, the objective can be rearranged to a combination of linear functions:
@@
Noted that the objective function $T$ is not linear because $W=\sum_{m\in\mathcal{M}}w_{m}$ is the sum of decision variables, and the term $\frac{1}{W}$ introduces a nonlinear dependency, which makes the problem a nonlinear optimization problem.
Now, we put it all together to present a final model:
@@
Constraint (23) requires that the window size $w_{n}$ must be a positive integer, the number of GPU layers $n_{m}$ must be a non-negative integer, and $n_{m}$ cannot exceed $w_{m}$. Constraint (24) requires that all devices are assigned an equal number of windows and all the windows are filled. This is not mandatory in our implementation, but can simplify the problem model. Constraint (28-30) ensure that devices categorized into sets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ meet the memory condition outlined in Cases 1-3. Similarly, Constraints (31-33) ensure that devices assigned to set $\mathcal{M}_{4}$ meet the memory condition outlined in Case 4. $b_{m}^{cio}$ in eq. (34) is a platform-independent constant. Constraints (35-36) ensure that the VRAM used by CUDA or the shared memory used by Metal does not exceed the available capacity. Here, $d_{m,\text{cuda}}^{\text{avail}}$ denotes the available GPU private memory for CUDA, and $d_{m,\text{metal}}^{\text{avail}}$ denotes the maximum working set size recommended by Metal. Note that the output layer weights (of $b_{o}$ bytes) are kept in Metal’s shared memory but run on the CPU by default.
This is an integer linear fractional programming (ILFP) problem because the numerator is a linear function of the decision variables $w_{m},n_{w}$ and the denominator $W$ is also a linear function of $w_{m}$. Moreover, the constraints are linear inequalities. Even when indicator variables $\mathbb{I}_{\text{macOS}},\mathbb{I}_{\text{Linux}},\mathbb{I}_{\text{Android}},\mathbb{I}_{\text{cuda}},\mathbb{I}_{\text{metal}},\mathbb{I}_{m=1}$ appear, each device’s platform is known in advance, so these indicators are fixed, they just activate or deactivate certain linear constraints for each device.
Next, we transform the above model into its vectorized form. Let the decision variables be ${\bm{w}}^{\mathrm{T}}=[w_{1},w_{2},\cdots,w_{M}]$, ${\bm{n}}^{\mathrm{T}}=[n_{1},n_{2},\cdots,n_{M}]$, and the coefficients ${\bm{a}},{\bm{b}},{\bm{c}}$ be:

### v2

@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
For the device loading time, processing cores must load data from RAM/VRAM into registers before executing instructions, which incurs latency. However, the theoretical RAM/VRAM bandwidth does not determine this latency because applications often fail to fully utilize the bandwidth, and multi-level caching also has a significant influence. To capture these effects, our device profiler implements an operator to read data from RAM/VRAM into registers. By measuring its latency at data volumes similar to the tensor sizes, we obtain practical throughputs $\{\mathcal{T}_{m}^{\text{cpu}},\mathcal{T}_{m}^{\text{cuda}},\mathcal{T}_{m}^{\text{metal}}\}$. Next, we calculate the data volume that needs to be loaded into registers during each token step, which typically consists of the weight data and the KV cache. In the model profiler, we record the total bytes of weight data for the input and output layers as $b_{i},b_{o}$, and for each layer as $b$. Additionally, the KV cache size for each layer is $2(h_{k}e_{k}+h_{v}e_{v})n_{kv}$, where $n_{kv}$ is the number of tokens for which the cache is stored. Then, the device loading time $T_{m}^{\text{dev\_load}}$ can be expressed as $T_{m}^{\text{dev\_load}}=(\frac{l_{m}-l_{m}^{\text{gpu}}}{\mathcal{T}_{m}^{\text{cpu}}}+\frac{l_{m}^{\text{gpu}}}{\mathcal{T}_{m}^{\text{gpu}}})(b+2(h_{k}e_{k}+h_{v}e_{v})n_{kv})+\frac{b_{i}/V+b_{o}}{\mathcal{T}_{m}^{\text{cpu}}}\cdot\mathbb{I}_{m=1}$, where $\mathcal{T}_{m}^{\text{gpu}}$ depends on the hardware: it equals $\mathcal{T}_{m}^{\text{metal}}$ for Metal and $\mathcal{T}_{m}^{\text{cuda}}$ for CUDA, and $V$ is the vocabulary size.
@@
Estimation of disk loading latency $T_{m}^{\text{disk}}$. Prima.cpp is designed to run on memory-constrained devices, so it cannot load the entire model into RAM. To address this, prima.cpp uses mmap to manage model weights. By using mmap, model weights are loaded into memory from disk only when needed for computation, and the OS will release inactive memory-mapped (mmapped) pages when memory pressure is high. This prevents OOM risks but incurs significant disk I/O latency, because evicted pages must be reloaded from disk the next time they are needed. To estimate this disk loading latency, it is necessary to determine the data volume that mmap needs to reload in each token step. This is a challenging task because different OSs exhibit very different memory management behaviors and high variability.
On macOS (without Metal) and Linux, the OS gradually reclaims memory. When memory pressure is moderate, i.e., when $b_{lio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}>d_{m}^{\text{avail}}$, mmapped pages are released incrementally until the pressure is alleviated. As a result, some weight data remain in the page cache, and the amount of data that mmap needs to reload is $\max(b_{lio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}},0)$, where $b_{lio}=(l_{m}-l_{m}^{\text{gpu}})b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}$, $2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})$ is the KV cache size, and $c^{\text{cpu}}$ is the compute buffer size. If CUDA is enabled on Linux, the model weights in private VRAM are locked by the CUDA driver, keeping them resident so no disk I/O occurs. Therefore, the disk loading latency for macOS (without Metal) and Linux can be estimated as
@@
For $T_{m,\text{macOS(no Metal)}}^{\text{disk}}$, $l_{m}^{\text{gpu}}=0$, and $s_{m}^{\text{disk}}$ is the random-read throughput of the disk. On Linux, mmap is configured for sequential access, so $s_{m}^{\text{disk}}$ is now the sequential-read throughput of the disk.
When Metal is enabled on macOS, the behavior changes. Metal loads mmapped pages into shared memory, and the OS prioritizes retaining these pages. That is, the OS is more inclined to swap out or compress active pages while keeping mmapped model weight pages in shared memory intact. However, when memory is exhausted (with free and inactive pages exhausted, the compression pool nearing saturation, and heavy swap usage), macOS will release these mmapped pages more aggressively. This may cause the entire model weights to be repeatedly reloaded and released. As a result, when the required memory exceeds the total available memory, i.e., when $l_{m}b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}l_{m}+c^{\text{cpu}}+c^{\text{gpu}}>d_{m,\text{metal}}^{\text{avail}}$, device $d_{m}$ needs to reload $l_{m}b+(b_{i}/V+b_{o})\cdot\mathbb{I}_{m=1}$ bytes in each token step. Here, $d_{m,\text{metal}}^{\text{avail}}$ denotes the maximum working set size recommended by Metal. By measuring the random-read throughput $s_{m}^{\text{disk}}$ of the disk, we can calculate the disk loading latency for macOS (with Metal) as:
@@
$+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}l_{m}+c^{\text{cpu}}+c^{\text{gpu}}-d_{m,\text{metal}}^{\text{avail}}\Big),\frac{b_{i}}{s_{m}^{\text{disk}}V}\Big).$ (16)
When running on Android devices, the OS prioritizes swapping out inactive pages to disk, such as memory used by background applications, to ensure that the active application runs smoothly. As a result, the available RAM for prima.cpp can be higher than expected because the OS swaps cold pages to disk, freeing some memory. Thus, on Android, the number of bytes that mmap needs to reload is $\max(b_{bio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}}-d_{m}^{\text{swapout}},0)$, where $d_{m}^{\text{swapout}}=\min(\max(0,b_{bio}+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}(l_{m}-l_{m}^{\text{gpu}})+c^{\text{cpu}}-d_{m}^{\text{avail}}),\min(d_{m}^{\text{bytes\_can\_swap}},d_{m}^{\text{swap\_avail}}))$ represents the data bytes that are swapped to disk, $d_{m}^{\text{bytes\_can\_swap}}$ is the data bytes of currently used memory that can be swapped out, and $d_{m}^{\text{swap\_avail}}$ is the total available swap space. Then we have:
@@
By aggregating them, we obtain a unified expression compatible with cross-platform devices:
[…]
@@
where $\mathbb{I}_{\text{macOS (no Metal)}},\mathbb{I}_{\text{macOS (with Metal)}},\mathbb{I}_{\text{Linux}},\mathbb{I}_{\text{Android}}$ are indicator functions. This expression can be easily extended to include new OSs.
Estimation of network communication latency $T_{m}^{\text{comm}}$. In prima.cpp, devices are connected in a ring, and each device receives inputs from its predecessor, processes them, and forwards the outputs to the next device. After a device completes the computation for one layer window, it transmits the result ($e$ values in the FP32 format, totaling $4e$ bytes) to the next device for further computation in the next layer window. During each token step, the number of network communications on device $d_{m}$ equals the number of layer windows, which is $\mathcal{W}_{m}=\left\lfloor\frac{L}{W}\right\rfloor+\min(1,\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$. By measuring the latency $t_{m}^{\text{comm}}$ of transmitting $4e$ bytes between adjacent devices, we can estimate the network communication latency on device $d_{m}$ as:
@@
By aggregating these latencies, our objective becomes:
@@
To remove the max operator, we decompose the disk loading latency into multiple terms, separately accounting for whether memory is sufficient or not. Let $\mathcal{M}$ be the set of all devices, and $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$ be the subsets of devices that satisfy the respective conditions in Cases 1-4, where $\mathcal{M}_{1}\cap\mathcal{M}_{2}\cap\mathcal{M}_{3}\cap\mathcal{M}_{4}=\emptyset$ and $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}\cup\mathcal{M}_{4}=\mathcal{M}$.
@@
Case 4 (OS with sufficient RAM or low disk speed): In these cases, the physical RAM is large enough to hold the model weights or the disk speed is slow (i.e., $s_{m}^{\text{disk}}<s_{\text{threshold}}^{\text{disk}}$). As a result, no disk loading is expected, except for the latency incurred during lookup table access, thus $T_{m}^{\text{disk}}=\frac{b_{i}}{s_{m}^{\text{disk}}V},m\in\mathcal{M}_{4}$.
@@
Let $\frac{L}{W}$ be an integer (i.e., $R=0$), where $W=\sum_{m\in\mathcal{M}}w_{m}$. Then all devices are assigned an equal number of windows, and all windows are filled.
Thus, we have $l_{m}=\frac{w_{m}L}{W}$, $l_{m}^{\text{gpu}}=\frac{n_{m}L}{W}$, $\mathcal{W}_{m}=\frac{L}{W}$. Let $b^{\prime}=b+2(h_{k}e_{k}+h_{v}e_{v})n_{kv}$, $\alpha_{m}=\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+t_{m}^{\text{kv\_cpy,cpu}}+\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{cpu}}}$, $\beta_{m}=\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{gpu},q}}-\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+t_{m}^{\text{kv\_cpy,gpu}}-t_{m}^{\text{kv\_cpy,cpu}}+\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{gpu}}}-\frac{b^{\prime}}{\mathcal{T}_{m}^{\text{cpu}}}$, $\xi_{m}=(t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}})(1-\mathbb{I}_{m}^{\text{UMA}})+t_{m}^{\text{comm}}$, $\kappa=\sum_{q\in\mathcal{Q}}\frac{f_{1,\text{out}}^{q}}{s_{1}^{\text{cpu},q}}+\frac{b_{i}/V+b_{o}}{\mathcal{T}_{1}^{\text{cpu}}}+\frac{b_{i}/V}{s_{1}^{\text{disk}}}+\frac{b_{o}}{s_{1}^{\text{disk}}}\cdot\mathbb{I}_{1\notin\mathcal{M}_{4}}+\sum_{m\in\mathcal{M}_{1}\cup\mathcal{M}_{3}}\frac{c^{\text{cpu}}-d_{m}^{\text{avail}}-d_{m}^{\text{swapout}}\cdot\mathbb{I}_{\text{Android}}}{s_{m}^{\text{disk}}}$, where $\alpha_{m},\beta_{m},\xi_{m}$ are platform-specific constants and $\kappa$ is a global constant. Then, we add the first general term to the three platform-specific terms and obtain:
@@
This objective is a sum over device sets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$. Each summand involves expressions linear in $w_{m}$ and $n_{m}$, plus platform-specific constant terms. To clarify the form, we define a linear function $f(a,b,c)=aw_{m}+bn_{m}+c$, where the platform-specific constants $a,b,c$ are independent of decision variables $w_{m},n_{m}$. Consequently, the objective can be rearranged to a combination of linear functions:
@@
Note that the objective $T$ is nonlinear because $W=\sum_{m\in\mathcal{M}}w_{m}$ depends on the decision variables, and the term $\frac{1}{W}$ introduces nonlinearity. Now, we put everything together:
@@
Constraint (22) requires that the window size $w_{m}$ must be a positive integer, the number of GPU layers $n_{m}$ must be a non-negative integer, and $n_{m}$ cannot exceed $w_{m}$. Constraint (23) requires that all devices be assigned an equal number of windows and all windows be filled. Constraints (27-29) ensure that devices categorized into sets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ meet the memory condition outlined in Cases 1-3. Similarly, Constraints (30-32) ensure that devices assigned to set $\mathcal{M}_{4}$ meet the memory condition outlined in Case 4. $b_{m}^{cio}$ in Eq. (33) is a platform-independent constant. Constraints (34-35) ensure that the VRAM used by CUDA or the shared memory used by Metal does not exceed the available capacity. Here, $d_{m,\text{cuda}}^{\text{avail}}$ denotes the available GPU private memory for CUDA, and $d_{m,\text{metal}}^{\text{avail}}$ denotes the maximum working set size recommended by Metal.
This is an integer linear fractional program (ILFP) because the numerator is a linear function of the decision variables $w_{m},n_{w}$, and the denominator $W$ is also a linear function of $w_{m}$. Moreover, the constraints are linear inequalities. The platform indicators $\mathbb{I}_{\text{macOS}},\mathbb{I}_{\text{Linux}},\mathbb{I}_{\text{Android}},\mathbb{I}_{\text{cuda}},\mathbb{I}_{\text{metal}},\mathbb{I}_{m=1}$ are known a priori, and they activate/deactivate corresponding linear constraints for each device.
Next, we transform the model into a vectorized form. Let the decision variables be ${\bm{w}}^{\mathrm{T}}=[w_{1},w_{2},\cdots,w_{M}]$, ${\bm{n}}^{\mathrm{T}}=[n_{1},n_{2},\cdots,n_{M}]$, and the coefficients ${\bm{a}},{\bm{b}},{\bm{c}}$ be: